OpenStudy (anonymous):

Find the volume of an ellipsoid with a semi-major axis of length 10 and semi-minor axes of lengths 6 and 4. Use 3.14 for pi and round your answer to the nearest tenth.

OpenStudy (anonymous):

1,004.8 320 883.2--- my answer 925.3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large V = \frac{4}{3}*\pi*a*b*c\]

jimthompson5910 (jim_thompson5910):

In this case, a = 10 b = 6 c = 4

OpenStudy (anonymous):

320

jimthompson5910 (jim_thompson5910):

don't forget about pi

jimthompson5910 (jim_thompson5910):

\[\Large V = 320\pi \approx \ ???\]

OpenStudy (anonymous):

1,004.8

jimthompson5910 (jim_thompson5910):

you got it